# Learning the Mechanism of Catastrophic Forgetting: A Perspective from Gradient Similarity

### v1

@@ Abstract
Catastrophic forgetting during knowledge injection severely undermines the continual learning capability of large language models (LLMs). While existing methods attempt to mitigate this issue, they often lack a foundational theoretical explanation. In this paper, we establish a theoretical framework from the perspective of gradient to explain catastrophic forgetting. First, we theoretically prove that strongly negative gradient similarity is the fundamental cause of catastrophic forgetting. Subsequently, gradient similarity is employed to identify two types of neurons during catastrophic forgetting: conflicting neurons, which induce forgetting and account for 50%–75% of all neurons, and collaborative neurons, which mitigate forgetting and account for 25%-50%. Ultimately, a novel knowledge injection method named Collaborative Neural Learning (CNL) is proposed. By freezing conflicting neurons and only updating collaborative neurons, CNL theoretically eliminates catastrophic forgetting under the assumption of an infinitesimal learning rate $\eta$ and exactly known Mastered Set. Studies on five LLMs, four datasets, and four optimizers prove that CNL achieves zero forgetting under in-set experiments, and reduces the level of catastrophic forgetting by 59.1%–81.7% under out-of-set experiments.
@@ 1 Introduction
Large language models (LLMs) have shown exceptional knowledge capacity and achieved remarkable potential across a broad range of domains Achiam et al. (2023); Touvron et al. (2023). To continuously expand their knowledge breadth, various knowledge injection methods have been proposed Lewis et al. (2020); Wang et al. (2024); Ke et al. (2023). Although these methods enable LLMs to acquire new knowledge, they also cause the forgetting of previously mastered knowledge Zhou et al. (2025). This phenomenon, known as catastrophic forgetting Korbak et al. (2022), severely hinders the continual improvement of LLMs Zhou et al. (2024). To mitigate catastrophic forgetting, it is crucial to understand its underlying mechanisms.
Recent studies have attempted to analyze catastrophic forgetting from a “similarity” perspective Bai et al. (2024); Wu et al. (2024). Specifically, they measure the similarity between injected knowledge and previously mastered knowledge using embedding representations Mitchell et al. (2022) and knowledge graphs Tan et al. (2025); Zhu et al. (2025). Results show that knowledge highly similar to injected knowledge is more likely to be forgotten Luo et al. (2025); Li et al. (2024). While these studies provide valuable insights into catastrophic forgetting, they fail to rigorously and theoretically reveal its underlying mechanisms.
Additionally, numerous approaches have been proposed to mitigate catastrophic forgetting, including replay based Huang et al. (2024), parameter-efficient fine-tuning based Li et al. (2025), and knowledge editing based approaches Ge et al. (2024). While they alleviate forgetting to some extent, they lack a rigorous theoretical foundation and therefore cannot avoid catastrophic forgetting at theoretically.
Our work explores the mechanisms of catastrophic forgetting through both theoretical analysis and experimental validation. More specifically, we seek to answer the following three key questions regarding catastrophic forgetting:
$\bullet$ What conditions lead to catastrophic forgetting? In Section 3, we demonstrate through rigorous theoretical derivation and extensive empirical results that strongly negativee gradient similarity between injected knowledge and mastered knowledge underlies catastrophic forgetting.
@@
$\bullet$ Which neurons are responsible for catastrophic forgetting? In Section 4, we apply gradient similarity to reveal two distinct types of neurons: conflicting neurons and collaborative neurons. Conflicting neurons induce forgetting and comprise 50%–75% of all neurons, whereas collaborative neurons alleviate forgetting and constitute 25%–50%.
$\bullet$ How to prevent catastrophic forgetting? In Section 5, we propose a novel knowledge injection method, CNL, which freezes conflicting neurons and trains collaborative neurons. CNL theoretically prevent forgetting under the assumption of an infinitesimal learning rate $\eta$ and exactly known Mastered Set. Experiments on five models, four datasets, and four optimizers demonstrate CNL effectively mitigates forgetting under both in-set and out-of-set conditions.
To ensure the reproducibility of this study, all of the codes and datasets are included in Supplementary Material. Code and datasets are available at: https://github.com/yangmutian/Collaborative-Neuron-Learning
Figure 1: Conceptual illustration of our framework. We reveal the mechanism of catastrophic forgetting and propose a novel method to prevent its occurrence from a perspective of gradient similarity.
2 Definition of Learning and Forgetting from the Perspective of Gradient
@@ 2.1 Learning
The neural network of LLMs is defined as $f_{\theta}:\mathcal{X}\to\mathcal{Y}$, where the mapping from the input space to the output space is determined by parameters $\theta=(\theta_{1},\dots,\theta_{|\theta|})\in\mathbb{R}^{|\theta|}$.
Let $\mathcal{I}$ denote a set of knowledge samples to be injected
@@
Loss on a single injection sample $(x,y)\in\mathcal{I}$ is expressed as $L^{(x,y)}(\theta)=\ell(f_{\theta}(x),y)$, where $\ell(\cdot,\cdot)$ denotes the loss function. The total loss on Injection Set $\mathcal{I}$ is expressed as the mean loss over all samples
@@
Assuming an infinitesimal learning rate $\eta$, the loss after parameter update can be approximated using a first-order Taylor expansion
@@
Due to the non-negativity of the inner product, we have $\Delta\mathcal{L}^{\mathcal{I}}=-\eta S(\mathcal{I},\mathcal{I})\leq 0$. In practice, parameter updates are rarely zero strictly, leading to
$\Delta\mathcal{L}^{\mathcal{I}}=-\eta S(\mathcal{I},\mathcal{I})<0$ (7)
This derivation defines the process of learning: under a sufficiently small learning rate $\eta$, knowledge injection monotonically decreases the loss on Injection Set $\mathcal{I}$.
@@ 2.2 Forgetting
The loss on Mastered Set is defined as $\mathcal{L}^{\mathcal{M}}(\theta)$. After parameter update according to Eq. (3), the loss becomes $\mathcal{L}^{\mathcal{M}}\left(\theta-\eta\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta)\right)$. Assuming an infinitesimal learning rate $\eta$, the expression is expanded through first-order Taylor approximation:
@@
We define the Gradient Similarity between Mastered Set $\mathcal{M}$ and Injection Set $\mathcal{I}$ as
@@
Therefore, during knowledge injection, the loss change on Mastered Set is given by
@@
Eq. (12) reveals a critical insight: the loss on the Mastered Set increases when the gradient similarity is negative. Therefore, mastered samples with more negative gradient similarity (i.e., stronger alignment in the opposite direction) are more susceptible to catastrophic forgetting.
@@ 3 Gradient Similarity Measures Catastrophic Forgetting
To validate the conclusion of Eq. (12), experiments on five models and four datasets were conducted.
3.1 Experiment Procedure
Datasets: We employed standard benchmarks including Massive Multitask Language Understanding (MMLU) Hendrycks et al. (2020), Medical Question Answering (MedQA) Jin et al. (2021), AI2 Reasoning Challenge (ARC-C) Clark et al. (2018), and Commonsense Question Answering (CSQA) Talmor et al. (2019). The detailed information of datasets is shown in Appendix A.
Models: The experiments were conducted on Qwen Yang et al. (2024) and LLaMA Grattafiori et al. (2024) families, including Qwen2.5 1.5B, Qwen2.5 3B, Qwen2.5 7B, LLaMA3.2 1B, and LLaMA3.2 3B.
3.2 Catastrophic Forgetting Happens during Knowledge Injection
To distinguish Injection Set $\mathcal{I}$ and Mastered Set $\mathcal{M}$, LLMs were asked to answer the questions in datasets (inference details in Appendix B). Correctly answered questions were categorized into $\mathcal{M}$, while incorrect ones formed $\mathcal{I}$. Appendix C exhibits the distribution of Injection Set $\mathcal{I}$ and Mastered Set $\mathcal{M}$.
Subsequently, LLMs were trained on Injection Set $\mathcal{I}$ for knowledge injection (training details in Appendix  D). The correctly answered samples in Injection Set $\mathcal{I}$ were used to evaluate learning, while the incorrectly answered samples in Mastered Set $\mathcal{M}$ were used to measure forgetting. Results in Appendix E reveal that while LLMs successfully learn new knowledge from Injection Set $\mathcal{I}$, they simultaneously suffer forgetting on Mastered Set $\mathcal{M}$.
Notably, we extended our evaluation to parameter-efficient fine-tuning (PEFT) methods and observed that techniques like Low-Rank Adaptation (LoRA) Hu et al. (2022) could not mitigate this catastrophic forgetting. Detailed experimental settings and results are provided in Appendix F.
3.3 Strongly Negative Gradient Similarity Leads to Forgetting
Ultimately, gradient similarities between samples in Mastered Set $\mathcal{M}$ and Injection Set $\mathcal{I}$ were computed. In light of Eq. (12), samples with negative gradient similarity were ranked by the magnitude of their similarity. The top $1/3$ with the largest magnitudes were assigned to the “Sim” group, representing strongly negative similarity, while the bottom $1/3$ were assigned to the “Dissim” group.
Results in Table 1 indicate that catastrophic forgetting is predominantly concentrated in the “Sim” group. This empirical evidence provides strong support for Eq. (12), confirming that strongly negative gradient similarity drives catastrophic forgetting.
Table 1: Relationship between gradient similarity and catastrophic forgetting. Results are reported in the format of “Number of forgotten questions (Percentage)”. Samples with negative gradient similarity were ranked by magnitude: the top $1/3$ with the largest magnitudes were labeled “Sim” (High Negative Similarity), while the bottom $1/3$ were labeled “Dissim”. Results indicate that catastrophic forgetting occurs primarily in the “Sim” group.
@@ 3.3 Strongly Negative Gradient Similarity Leads to Forgetting
4 Conflicting Neurons Cause Catastrophic Forgetting
While LLMs exhibit overall forgetting on Mastered Set, it remains unclear whether forgetting is driven by all neurons or only a subset. Herein, gradient similarity is employed to identify the neurons contributing to forgetting. It should be clarified that in this paper, “neurons” and “parameters” carry essentially the same meaning. We use parameters when emphasizing their vector form for computation, and neurons when emphasizing their count.
The gradient similarity of the $j$-th neuron $\theta_{j}$ between Mastered Set $\mathcal{M}$ and Injection Set $\mathcal{I}$ is defined as
$\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{M}}(\theta)$ and $\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{I}}(\theta)$ represent the components of gradients on $\theta_{j}$ for Mastered and Injection Set, respectively.
The global gradient similarity of LLMs is equivalent to the sum of gradient similarities of individual neurons. Therefore, Eq. (11) is expressed as
$S(\mathcal{M},\mathcal{I})$ $=\nabla_{\theta}\mathcal{L}^{\mathcal{M}}(\theta)^{\top}\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta)$
$=\sum_{j=1}^{|\theta|}\left(\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{M}}(\theta)\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{I}}(\theta)\right)$
$=\sum_{j=1}^{|\theta|}s_{\theta_{j}}(\mathcal{M},\mathcal{I})$ (14)
According to Eq. (12), if the gradient similarity of neuron $\theta_{j}$ with respect to Mastered Set $\mathcal{M}$ and Injection Set $\mathcal{I}$ is negative, this neuron contributes to an increase in loss and induce forgetting. We refer to such neurons as Conflicting Neurons. The set of conflicting neurons is defined as
$\theta_{CF}:=\left\{\theta_{j}\mid s_{\theta_{j}}(\mathcal{M},\mathcal{I})<0\right\}$ (15)
Conversely, if the gradient similarity of neuron $\theta_{j}$ is positive, this neuron contributes to a decrease in loss and mitigate forgetting. Such neurons are refered as Collaborative Neurons. The set of collaborative neurons is defined as
$\theta_{CB}:=\left\{\theta_{j}\mid s_{\theta_{j}}(\mathcal{M},\mathcal{I})\geq 0\right\}$ (16)
Table 2 presents that nearly 25%-50% of neurons contribute to positive gradient similarity and are termed collaborative neurons, while 50%-75% of neurons contribute to negative gradient similarity and are termed conflicting neurons. These collaborative neurons are insufficient to counteract the effect of conflicting neurons, thereby leading to the phenomenon of catastrophic forgetting in LLMs.
This result reveals the coexistence of two types of neurons during catastrophic forgetting, which provide theoretical foundation for the following knowledge injection method.
Table 2: Distribution of collaborative neurons (COLL) and conflicting neurons (CONF). "PROP" represents the proportion of a specific neuron type relative to the total number of neurons. “GRAD” represents the gradient similarity sum for that neuron type. “TOTAL” represents the gradient similarity sum for all neurons. Typically, collaborative neurons account for 25%–50% of all neurons, while conflicting neurons account for 50%–75%.
@@ 4 Conflicting Neurons Cause Catastrophic Forgetting
5 Preventing Catastrophic Forgetting by Freezing Conflicting Neurons
Since conflicting neurons are responsible for catastrophic forgetting, a natural question arises: can catastrophic forgetting be avoided by freezing conflicting neurons and training only collaborative neurons? We refer to this learning paradigm as Collaborative Neuron Learning (CNL).
5.1 Establishment of CNL
Table 3: Effectiveness of knowledge injection using FT and CNL. Results are reported in the format of “Number of questions (Percentage)”. FT leads to the forgetting of previously mastered knowledge comparable in magnitude to the newly injected knowledge, whereas CNL induces negligible catastrophic forgetting.

### v2

# Collaborative Parameter Learning: Mitigating Forgetting via Parameter-Level Gradient Analysis

@@ Abstract
Catastrophic forgetting during knowledge injection impairs the ability of large language models to acquire new knowledge without overwriting previously mastered knowledge. Recent studies analyze forgetting from a gradient-similarity perspective and mitigate forgetting through vector projection. However, these methods primarily characterize gradient similarity at the aggregate direction level, leaving the parameter-wise contributions to forgetting underexplored. In this paper, we decompose gradient similarity into parameter-wise contributions and identify two types of parameters during forgetting: Conflicting Parameters, whose updates contribute to forgetting and typically account for 50%–75% of parameters, and Collaborative Parameters, whose updates mitigate forgetting and account for 25%–50%. Based on this analysis, we propose Collaborative Parameter Learning (CPL), a parameter-wise training rule that freezes Conflicting Parameters and updates only Collaborative Parameters. Experiments comparing CPL with seven baseline methods show that CPL learns 20.2%–48.2% more questions with negligible forgetting, while reducing peak VRAM by approximately 3 GB per billion model parameters and computation time by 16.5%. Extensive evaluations on parameter consumption, out-of-set generalization, cross-prompt generalization, multimodal tasks, open-ended question answering, and multilingual settings demonstrate that CPL effectively mitigates forgetting across diverse scenarios.
@@ 1 Introduction
Large language models (LLMs) have shown exceptional knowledge capacity and achieved remarkable potential across a broad range of domains [1, 2]. To expand their knowledge breadth, various knowledge injection methods have been proposed [3, 4, 5]. Although these methods enable LLMs to acquire new knowledge, they also cause the forgetting of previously mastered knowledge [6]. This phenomenon, known as catastrophic forgetting [7], severely undermines the preservation of previously acquired knowledge in LLMs [8].
Existing literatures on mitigating catastrophic forgetting are typically categorized into four primary paradigms [9]. Regularization-based methods penalize changes to parameters critical for previously mastered knowledge [10, 11], while parameter-isolation techniques alleviate interference by confining updates to dedicated modules or parameter subsets [12, 13]. Among these paradigms, gradient-based methods are particularly significant because they formalize forgetting as an optimization-level conflict between injected and mastered knowledge [14, 15]. By characterizing interference through gradient similarity, these methods mitigate forgetting by projecting the update direction to reduce detrimental interference with mastered knowledge [16, 17, 18].
However, most existing gradient-based methods treat gradient interference as a single, aggregate vector-level quantity [16, 17, 18], enforcing strict orthogonality between newly injected and previously mastered knowledge to mitigate forgetting. As a result, they overlook the parameter-wise contributions to forgetting and impose overly restrictive constraints on knowledge injection.
Therefore, critical questions arise: do all parameters contribute to forgetting, or only a subset of them? How to eliminate the conflicting components between injected and mastered knowledge while preserving the collaborative components, rather than strictly enforcing orthogonality? This work attempts to answer these questions as follows:
$\bullet$ How to characterize catastrophic forgetting from the perspective of gradients? In Section 3, we analyze the relationship between gradient similarity and forgetting, revealing that samples with strongly negative gradient similarity are more susceptible to forgetting.
@@
$\bullet$ Which parameters are responsible for catastrophic forgetting? In Section 4, we reveal two distinct types of parameters: Conflicting Parameters and Collaborative Parameters. Conflicting Parameters induce forgetting and typically comprise 50%–75% of all parameters, whereas Collaborative Parameters alleviate forgetting and constitute 25%–50%.
$\bullet$ How to mitigate catastrophic forgetting? In Section 5, we propose a novel knowledge injection method, Collaborative Parameter Learning (CPL), which freezes Conflicting Parameters and trains Collaborative Parameters. By preserving the collaborative directions between injected and mastered knowledge, rather than enforcing strict orthogonality, CPL learns 20.2%–48.2% more questions with negligible forgetting. By replacing high-cost gradient projection with low-cost sign comparison, CPL reduces peak VRAM by 3 GB per billion model parameters and computation time by 16.5%.
$\bullet$ Is CPL robustness to diverse scenarios? In Section 6, we extend CPL to out-of-set generalization, cross-prompt transfer, multimodal tasks, open-ended question answering, and multilingual settings. Results indicate that CPL effectively mitigates forgetting in those settings.
The appendix provides additional experimental details and results.
Figure 1: Conceptual illustration of our framework. We decompose gradient similarity into parameter-wise contributions, and propose a parameter-wise training rule to mitigate forgetting.
2 Definition of Learning and Forgetting
@@ 2.1 Learning
The parameters of LLMs are denoted as $\theta=(\theta_{1},\dots,\theta_{|\theta|})\in\mathbb{R}^{|\theta|}$, where $\theta_{j}$ represents the $j$-th parameter. Let $\mathcal{I}$ denote a set of knowledge samples to be injected
@@
Loss on a single injection sample $(x,y)\in\mathcal{I}$ is expressed as $L^{(x,y)}(\theta)=\ell(f_{\theta}(x),y)$, where $\ell(\cdot,\cdot)$ denotes the loss function. The total loss on Injection Set $\mathcal{I}$ is expressed as
@@
The loss after parameter update is approximated using a first-order Taylor expansion
@@
Due to the non-negativity of the inner product, we have $\Delta\mathcal{L}^{\mathcal{I}}=-\eta S(\mathcal{I},\mathcal{I})\leq 0$. This derivation characterizes knowledge injection: under a sufficiently small learning rate, updating along the negative gradient of $\mathcal{I}$ decreases its loss in the first-order approximation.
@@ 2.2 Forgetting
The loss on $\mathcal{M}$ is defined as $\mathcal{L}^{\mathcal{M}}(\theta)$. After parameter update , the loss becomes $\mathcal{L}^{\mathcal{M}}\left(\theta-\eta\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta)\right)$. The expression is expanded through first-order Taylor approximation:
@@
We define the gradient similarity between $\mathcal{M}$ and $\mathcal{I}$ as
@@
Therefore, during knowledge injection, the loss change on $\mathcal{M}$ is given by
@@
Eq. (11) reveals a critical insight: the loss on $\mathcal{M}$ increases when the gradient similarity is negative. Therefore, mastered samples with more negative gradient similarity (i.e., stronger alignment in the opposite direction) are more susceptible to catastrophic forgetting.
@@ 3 Gradient Similarity Measures Catastrophic Forgetting
Table 1: Relationship between gradient similarity and forgetting. Results are reported in the format of “Number of forgotten questions (Percentage)”. Samples with negative gradient similarity are ranked by magnitude: the top $1/3$ with the largest magnitudes are labeled “Sim”, while the bottom $1/3$ are labeled “Dissim”. Results indicate that forgetting occurs primarily in the “Sim” group.
@@
3.1 Experiment Procedure
Datasets: We employ standard benchmarks including Massive Multitask Language Understanding (MMLU) [19], Medical Question Answering (MedQA) [20], AI2 Reasoning Challenge (ARC-C) [21], Commonsense Question Answering (CSQA) [22]. Additionally, Yahoo Answers [23], AG News [23], and DBPedia [23], NumGLUE-cm [24], NumGLUE-ds [24], C-STANCE [25], and MMBench [26] are used for extensive study.
Models: The experiments are conducted on Qwen [27] and LLaMA [28] families, including Qwen2.5 1.5B, Qwen2.5 3B, Qwen2.5 7B, LLaMA3.2 1B, and LLaMA3.2 3B. The multimodal evaluation is conducted on Qwen2.5vl 3B.
3.2 Catastrophic Forgetting Happens during Knowledge Injection
To distinguish Injection Set $\mathcal{I}$ and Mastered Set $\mathcal{M}$, LLMs are asked to answer the questions in datasets (inference details in Appendix B). Correctly answered questions are grouped into $\mathcal{M}$, while incorrect ones form $\mathcal{I}$. Appendix C exhibits the distribution of $\mathcal{I}$ and $\mathcal{M}$.
LLMs are trained on $\mathcal{I}$ as detailed in Appendix  D. Samples in $\mathcal{I}$ are used to evaluate learning, while samples in $\mathcal{M}$ are used to measure forgetting. Table 8 reveals that while LLMs successfully learn new knowledge from $\mathcal{I}$, they simultaneously suffer forgetting on $\mathcal{M}$. Similar forgetting is also observed in Low-Rank Adaptation (LoRA) setting [13] as shown in Appendix F.
3.3 Strongly Negative Gradient Similarity Leads to Forgetting
Gradient similarity between samples in $\mathcal{M}$ and $\mathcal{I}$ is computed. In light of Eq. (11), samples with negative gradient similarity are ranked by the magnitude of their similarity. The top $1/3$ with the largest magnitudes are assigned to the “Sim” group, representing strongly negative similarity, while the bottom $1/3$ are assigned to the “Dissim” group.
Table 1 indicates that forgetting is predominantly concentrated in the “Sim” group. This evidence provides support for Eq. (11), confirming that strongly negative gradient similarity drives forgetting.
4 Conflicting Parameters Cause Catastrophic Forgetting
Table 2: Distribution of Collaborative Parameters (COLL) and Conflicting Parameters (CONF). “PROP” represents the proportion of a specific parameter. “GRAD” represents the gradient similarity sum. “TOTAL” represents the gradient similarity sum for all parameters. Collaborative Parameters typically account for 25%–50%, while Conflicting Parameters account for 50%–75%.
@@ 4 Conflicting Parameters Cause Catastrophic Forgetting
While LLMs exhibit overall forgetting on $\mathcal{M}$, it remains unclear whether forgetting is driven by all parameters or only a subset. To answer this question, gradient similarity of the $j$-th parameter between $\mathcal{M}$ and $\mathcal{I}$ is defined as
$\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{M}}(\theta)$ and $\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{I}}(\theta)$ represent the components of gradients on $\theta_{j}$ for $\mathcal{M}$ and $\mathcal{I}$, respectively. The global gradient similarity of LLMs is equivalent to the sum of gradient similarities of individual parameters. Therefore, Eq. (10) is expressed as
$S(\mathcal{M},\mathcal{I})$ $=\sum_{j=1}^{|\theta|}\left(\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{M}}(\theta)\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{I}}(\theta)\right)$
$=\sum_{j=1}^{|\theta|}s_{\theta_{j}}(\mathcal{M},\mathcal{I})$ (13)
If the gradient similarity of $\theta_{j}$ with respect to $\mathcal{M}$ and $\mathcal{I}$ is negative, this parameter contributes to an increase in loss and induces forgetting. We refer to such parameters as Conflicting Parameters. Conversely, if the gradient similarity of $\theta_{j}$ is positive, this parameter contributes to a decrease in loss and mitigates forgetting. Such parameters are referred to as Collaborative Parameters.
Table 2 describes the distribution of parameters. Approximately 25%–50% of parameters have positive gradient similarity and are termed Collaborative Parameters, while 50%–75% have negative contributions and are termed Conflicting Parameters. Moreover, the positive contributions from Collaborative Parameters are generally insufficient to offset the stronger negative contributions from Conflicting Parameters, causing the overall gradient similarity to remain negative.
5 Mitigating Catastrophic Forgetting by Freezing Conflicting Parameters
5.1 Establishment of CPL
Table 3: Performance of knowledge injection using FT and CPL. Results are reported in the format of “Number of questions (Percentage)”. FT leads to significant forgetting of previously mastered knowledge, whereas CPL induces negligible catastrophic forgetting when $\mathcal{M}$ is fully available.